# Dynamic Language Group-Based MoE: Enhancing Efficiency and Flexibility for Code-Switching Speech Recognition

### v1

@@ 3.3 Analysis of expert capacity and top-k
Baseline-18 173M 173M(top1) 6.65 4.50 24.25
DLG-MoE-8e 228M 100M(top1) 6.16 4.12 22.88
118M(top2) 6.11 4.08 22.78
3.4 Streaming Capability
@@
Mode Specify Lang. EN ZH CS
NonStreaming 5.29 2.42 8.92
Streaming(640ms) 6.21 2.82 9.89
Streaming(320ms) 6.69 3.02 10.31
Streaming(320ms) 6.64 2.98 10.31
In this section, we tested the model’s streaming recognition ability with different chunk sizes. As illustrated in Table 3, our model supports inference with varying chunk sizes while maintaining reasonable performance in streaming scenarios. The method also allows specifying a language, sending all frames to the designated language expert group to avoid incorrect routing during streaming. Although we did not observe significant performance improvements, this might be more noticeable on out-of-domain test sets. Fig.2 shows the visualization of the routing results of the language router in $S11$. We found that the language router has high routing accuracy, especially for monolingual utterances. It also tends to dispatch the first few frames of English to the Zh-Group, which may correspond to frames with no speech. For CS utterances, a high routing accuracy was also observed.
Fig. 2: Visualisation of SLR’s frame-level routing results, where red represents ‘Zh’ and green represents ‘En’.
4 Conclusions
This paper presents an efficient and flexible MoE structure. Through the proposed language-group-based routing and dynamic top-$k$ mechanism, the model achieves SOTA performance, significantly improving over baselines with comparable computational costs. The study also investigates the impact of expert capacity and different top-$k$ strategies on model performance, demonstrating that language groups and the proposed dynamic top-$k$ strategy are beneficial for flexibly expanding expert capacity. In the future, we will explore the potential of the proposed model across more languages.

### v2

@@ 3.3 Analysis of Expert Capacities and Top-k
Baseline-18 173M 173M(top-1) 6.65 4.50 24.25
DLG-MoE-8e 228M 100M(top-1) 6.16 4.12 22.88
118M(top-2) 6.11 4.08 22.78
3.4 Streaming Capability and Flexibility
@@
Mode Size Specify Lang. EN ZH CS
NonStreaming 103M 5.29 2.42 8.92
Streaming(640ms) 103M 6.21 2.82 9.89
Streaming(320ms) 103M 6.69 3.02 10.31
Streaming(320ms) 72M 6.64 2.98 -
[…]
Fig. 2: Visualisation of SLR’s frame-wise routing results, where red represents Chinese and green represents English.
4 Conclusions and future work
This paper proposes an efficient MoE model to address the CS scenario. Through the proposed hierarchical routing and dynamic language groups, the model achieves SOTA performance, significantly improving over baselines with comparable computational costs. The study also investigates the impact of expert capacity and different top-$k$ strategies on model performance, demonstrating that language groups and the proposed dynamic top-$k$ strategy are beneficial for flexibly expanding expert capacity. In the future, we will explore the potential of the proposed model across more languages.